31 A continuous random variable X has the following probability density function:
f(x) = 1/4, 0 ≤ x ≤ 4
Find the following probabilities:

a. P(X ≤ 1)
b. P(X ≤ 2)
c. P(1 ≤X ≤ 2)
d. P(X = 3)
I am totally lost.

Answer by stanbon(75887)

A continuous random variable X has the following probability density function:
f(x) = 1/4, 0 ≤ x ≤ 4'
Find the following probabilities:
Sketch the density function as a rectangle with base = 4 and height = (1/4)
a. P(X ≤ 1) = 1*(1/4) = 1/4
-------------------------------------------
b. P(X ≤ 2) = 2*(1/4) = 1/2
-------------------------------------------
c. P(1 ≤X ≤ 2) = (2-1)*(1/4) = 1*(1/4) = 1/4
-------------------------------------------
d. P(X = 3) = 3(1/4) = 3/4
--------------
Cheers,
Stan H.
-----------------